# Graph Neural Networks for Job Shop Scheduling Problems: A Survey

### v1

@@ 5 Discussion
Regarding FSPs: 1) The GNN applications in FSPs are not as popular as in JSSPs, though these two types of problems are very similar. 2) The GNN-based methods use different types of graphs to represent FSPs, such as bipartite, complete, and multiple graphs. There is a lack of comparative studies to determine the most preferred graph type. 3) The existing works on FSPs generally employ the commonly used GNN types, such as GCN, GAT, and GIN, to capture the embeddings of jobs or machines. Additional MLPs take in the embeddings for subsequent job or machine selection. A special case is that GCN can also be used to learn the representation of the Gantt chart. 4) Similar to JSSPs, most works on FSPs still used REINFORCE or PPO algorithms to train the neural networks. 5) The randomly generated data and public benchmark datasets such as TA and VRF (Vallada et al., 2015; Fernandez-Viagas and Framinan, 2020) are often tested in the experiments, and NEH is often taken as the main baseline. However, a comprehensive comparison of GNN-based methods is still missing.
@@ 6 Potential Research Directions
Performance: Despite advancements, the performance of existing GNN-based methods for static JSSPs is still far from optimal and outperformed by the top-performing heuristic or meta-heuristic methods. Hence, designing powerful learning-based algorithms to surpass state-of-the-art heuristics is still an open research problem. This goal can be approached from two main angles. Firstly, developing tailored GNNs that excel in learning effective policies for shop scheduling problems is crucial for improving performance. However, an optimal model cannot be achieved without employing a more effective problem representation. Therefore, exploring enhanced representation learning techniques alongside problem formulation strategies is recommended. Secondly, most existing GNN-based methods attempt to learn dispatch rules. Extending GNNs to work with more advanced solvers (e.g., meta-heuristic methods or exact solvers) is promising to further boost the performance. Thirdly, there is a need to explore more advanced DRL algorithms or machine learning paradigms to enhance performance further or improve generalization. For example, more advanced DRL algorithms should be investigated to further narrow the optimal gap for JSSPs. To improve generalization, meta-learning (Huisman et al., 2021) is a promising paradigm and has already shown effectiveness in the routing domain (Zhou et al., 2023a). Its combination with scheduling is a promising direction for future research.

### v2

@@ 5.2 Problem-Specific Observations:
  • 4.
    Regarding FSPs: 1) The GNN applications in FSPs are not as popular as in JSSPs. Compared to FSPs, JSSPs are featured by an additional precedence constraint, which requires that an operation cannot begin until the preceding operation of the same job is completed. This constraint enables JSSPs to encompass a wider array of real-case scenarios, highlighting the unique challenges and opportunities presented by JSSPs. 2) The GNN-based methods use different types of graphs to represent FSPs, such as bipartite, complete, and multiple graphs. There is a lack of comparative studies to determine the most preferred graph type. 3) The existing works on FSPs generally employ the commonly used GNN types, such as GCN, GAT, and GIN, to capture the embeddings of jobs or machines. Additional MLPs take in the embeddings for subsequent job or machine selection. A special case is that GCN can also be used to learn the representation of the Gantt chart. 4) Similar to JSSPs, most works on FSPs still used REINFORCE or PPO algorithms to train the neural networks. 5) The randomly generated data and public benchmark datasets such as TA and VRF (Vallada et al., 2015; Fernandez-Viagas and Framinan, 2020) are often tested in the experiments, and NEH is often taken as the main baseline. However, a comprehensive comparison of GNN-based methods is still missing.
According to the observations, most differences in approaches for each problem variant lie in graph representations, GNNs, and DRL algorithms. There is no one single technique that is best for all problem variants. In general, specialized techniques for each specific variant can bring benefits in delivering more effective solutions. This survey aims to summarize existing methods in JSSPs and present their differences for readers’ reference. We also highlighted more comprehensive and fair benchmarking is important to figure out the state-of-the-art techniques for each problem variant, so as to provide more advice and guidance for algorithm development.
@@ 6 Potential Research Directions
  • 2.
    Performance: Despite advancements, the performance of existing GNN-based methods for static JSSPs is still far from optimal and outperformed by the top-performing heuristic or meta-heuristic methods. Hence, designing powerful learning-based algorithms to surpass state-of-the-art heuristics is still an open research problem. This goal can be approached from two main angles. Existing methods often execute the policy learning in a constructive or improving heuristic for JSSPs, with the policy parameterized by GNNs and trained by DRL algorithms. In general, the GNNs’ role is to effectively capture and encode the complex, high-dimensional state space of the JSSP into low-dimensional embeddings that can be used to approximate a (favorable) rule in the heuristic (e.g., the dispatching rule in the constructive heuristic). GNNs also help in learning the interdependencies and constraints of the problem, thereby enhancing the decision-making capabilities of heuristics. In contrast, the role of RL is modeling the interactions between the GNN (i.e., agent) and the heuristic (i.e., environment) as a MDP process and updating the GNN to optimize the objective. Accordingly, we can improve the current learning-based methods from three main angles. Firstly, developing tailored GNNs that excel in learning effective policies for shop scheduling problems is crucial for improving performance. However, an optimal model cannot be achieved without employing a more effective problem representation. Therefore, exploring enhanced representation learning techniques alongside problem formulation strategies is recommended. Secondly, most existing GNN-based methods attempt to learn dispatch rules. Extending GNNs to work with more advanced solvers (e.g., meta-heuristic methods or exact solvers) is promising to further boost the performance. Thirdly, there is a need to explore more advanced DRL algorithms or machine learning paradigms to enhance performance further or improve generalization. For example, more advanced DRL algorithms should be investigated to further narrow the optimal gap for JSSPs. To improve generalization, meta-learning (Huisman et al., 2021) is a promising paradigm and has already shown effectiveness in the routing domain (Zhou et al., 2023a). Its combination with scheduling is a promising direction for future research.
  • 3.
@@
  • 6.
    Benchmarking: Benchmarking is essential for evaluating the effectiveness and efficiency of scheduling algorithms. Despite the availability of benchmarks for classical scheduling problems, there is a notable lack of benchmarks for many variants, especially for dynamic job shop scheduling problems (DyJSSPs) and problems with practical constraints. This gap hinders comprehensive evaluation and fair comparison of algorithms under consistent conditions. Currently, we are aware of only one benchmarking environment, Reijnen et al. (2023), that supports the evaluation of scheduling algorithms under static and dynamic conditions. Future research should focus on developing more unified benchmark instances and environments, enabling fair comparisons of proposed methods.